# Equation of the Perfect Fluid in the FRW Universe

### v1

@@ I Introduction
and the horizon radius is related with the thermodynamic volume, i.e. $V=V(r_{h})$, e.g. $V=4\pi r_{h}^{3}/3$ for a 4-dimensional spherically symmetric black holes. Therefore, the Hawking temperature $T$ can be regarded as a function of the thermodynamic volume $V$ and the thermodynamic pressure $P$ as well as other variables, i.e. T=T(V,P,)T=T(V,P,...), which can also be written as P=P(V,T,)P=P(V,T,...) and is a vdW-like equation of state. It has been found that for many black holes in AdS and dS spacetime, their equations of state have vdW-like critical points in the $P$-$V$ diagram that satisfy
@@
and there are phase transitions below the critical temperature [^1] The first example is the charged black hole in anti-de Sitter spacetimeKubiznak:2012wp .. Below the critical temperature, there are two phases or states-the small state and the large state similar to the liquid state and gas state of the vdW system, and they can co-exist at certain temperature and pressure with the same free energy. Apart from the co-existence states, the free energy of the two states are different, and the state with a lower free energy is thermodynamically stabler, so there will be a phase transition between the two states as pressure or temperature moves across the co-existence value. This is a first order phase transition as the first order derivative of the free energy is not continuous at this point although the free energy itself is continuous. Above the critical temperature, the two phases cannot be distinguished, i.e. there is only one phase, so the free energy and its arbitrary order derivatives are always continuous thus there is no phase transition. Interestingly, at the critical point, the free energy and its first order derivative are continuous but its second order derivative is not continuous, so there is a second order phase transition. Near the critical point, one can also define critical exponentsPelissetto:2000ek , which were found to be the same with those in the vdW system and the mean field theory in nearly all the cases, so they satisfy the scaling laws. Recently, critical exponents beyond mean field theory have been found and even violate scaling lawsHu:2024ldp ; Kong:2025lqz . People also found that black holes can undergo Joule-Thomson expansionOkcu:2016tgt , have reentrant phase transitionAltamirano:2013ane and triple pointAltamirano:2013uqa , etc.
Given the thermodynamic similarities between the F(L)RW(Friedmann-(Lemaître)-Robertson-Walker) universe and black holes, one expects the FRW universe to also exhibit a vdW-like equation of state [^2] It should be noted that in the cosmology community, ‘equation of state’ is usually referred to the relation between pressure $p$ and energy density $\rho$, i.e. $p=\omega\rho$, which is different from the vdW-like equation. In this paper, we use ‘equation of state’ to refer to the vdW-like equation as is often done in the study of black hole phase transitions.. The thermodynamic pressure of the FRW universe is identified with the work densityHayward:1997jp ; Cai:2005ra ; Akbar:2006kj ; Cai:2006rs ; Abdusattar:2021wfv ; Kong:2021dqd ; Kong:2022xny ; Abdusattar:2023hlj ; Chu:2025zuz ,
@@
where $\rho$ and $p$ are the energy density and pressure of the perfect fluid respectively. It can be written as an equation of state $P=P(V,T)$, where $V$ is the thermodynamic volume and $T$ is the Hawking temperature of the FRW universe. The equation of state depends on the theory of gravity and the dimension of the spacetime. For the (3+1)-dimensional FRW universe, it has been obtained in Einstein gravityAbdusattar:2021wfv , scalar-tensor theoriesKong:2021dqd ; Abdusattar:2023hlj , the brane world scenarioKong:2022xny , and quasi-topological gravityChu:2025zuz , etc. It has also been obtained for the high dimensional FRW universe in Einstein-Gauss-Bonnet gravitySaavedra:2023lds  [^3] I thank Haximjan Abdusattar for drawing my attention to this paper.. In many cases, the vdW-like phase transitions are found, but they occur above the critical temperature, which is different from the black hole cases.
In this paper, we are interested in the behavior of the perfect fluid in the FRW universe. The perfect fluid in the FRW universe also has a pressure as well as an energy density. If the perfect fluid is in an equilibrium with the Hawking radiation from the apparent horizon of the FRW universe, its temperature is also equal to the Hawking temperature. From the corresponding Friedmann’s equations, one can also get the equation of state for the perfect fluid in the FRW universe. The dimension of spacetime also affects the thermodynamic behaviors of the perfect fluid, so we study the perfect fluid in the general $D$-dimensional(or $d+1$ dimensional) FRW universe instead of the $4$-dimensional(or $3+1$ dimensional) FRW universe. In this paper, we only consider three common tensor theories of gravity, i.e. Einstein gravity, Gauss-Bonnet gravity and Lovelock gravityLovelock:1971yv ; Deruelle:1989fj ; Padmanabhan:2013xyr , and obtain the corresponding equation of state for the perfect fluid in the $D$-dimensional FRW universe. In some cases, the vdW-like phase transitions are also found, which depend on the theory of gravity and dimension of the FRW universe as expected. Phase transitions of the perfect fluid also occur above the critical temperature.
@@ II A Brief Introduction of the $D$-dimensional FRW Universe
where $a(t)$ is the time-dependent scale factor [^4] In the following, we may use $a$ instead of $a(t)$ for convenience., $k$ is the spatial curvature (with +1,0,1+1,0,-1 corresponding to spatial closed, flat and open respectively), and $\mathrm{d}\Omega_{n-1}^{2}$ is the metric of the unit (n1)(n-1)-dimensional sphere. One can also rewrite the metric as
@@ III Equation of the Perfect Fluid in Einstein Gravity
The equation of state can be regarded as a function $p=p(R_{A},T)$, in order to show its properties, one can either keep $R_{A}$ or $T$ fixed mathematically. However, for dynamical spacetime like the FRW universe, $R_{A}$ is not fixed, so we only fix $T$ and study the behavior of the isothermal lines, which shows the relation between the pressure and the radius of the apparent horizon. If $n=3$, for the isothermal line with temperature $T$, the pressure has a minimum $p_{min}=-2\pi T^{2}$ at $R=1/(4\pi T)$, and becomes zero at $R=1/(8\pi T)$, see FIG.1.666Here the temperature $T$ is given as 0.11, 0.1, 0.09 without unit, which seems strange for some readers. It should be noted that the figures are drawn to show the shape of the isothermal lines and the variation trend of the pressure with the temperature, so the specific value or unit is not so important but their comparisons. It is also a frequently used method in the references about black hole phase transitions, e.g.Kubiznak:2012wp . The values of the temperature in the following figures are chosen on the same spirit.
@@ VI Conclusions and Discussions
The results show that critical points and phase transitions of the perfect fluid depend on both the theories of gravity and the dimensions of the FRW universe. If the critical point exists, phase transition always takes place above the critical temperature, which appears to be a universal feature.

### v2

@@ I Introduction
and the horizon radius is related to the thermodynamic volume, i.e. $V=V(r_{h})$, e.g. $V=4\pi r_{h}^{3}/3$ for the 4-dimensional spherically symmetric black holes. Therefore, the Hawking temperature $T$ can be regarded as a function of the thermodynamic volume $V$ and the thermodynamic pressure $P$ as well as other variables, i.e. T=T(V,P,)T=T(V,P,...), which can also be written as P=P(V,T,)P=P(V,T,...) and is a vdW-like equation of state. It has been found that for many black holes in AdS and dS spacetime, their equations of state have vdW-like critical points in the $P$-$V$ diagram that satisfy
@@
and there are phase transitions below the critical temperature [^1] The first example is the charged black hole in anti-de Sitter spacetimeKubiznak:2012wp .. Below the critical temperature, there are two phases or states-the small state and the large state similar to the liquid state and gas state of the vdW system, and they can co-exist at certain temperature and pressure with the same free energy. Apart from the co-existence states, the free energy of the two states is different, and the state with a lower free energy is thermodynamically stabler, so there will be a phase transition between the two states as pressure or temperature moves across the co-existence value. This is a first order phase transition as the first order derivative of the free energy is not continuous at this point although the free energy itself is continuous. Above the critical temperature, the two phases cannot be distinguished, i.e. there is only one phase, so the free energy and its arbitrary order derivatives are always continuous, thus there is no phase transition. Interestingly, at the critical point, the free energy and its first order derivative are continuous but its second order derivative is not continuous, so there is a second order phase transition. Near the critical point, one can also define critical exponentsPelissetto:2000ek , which were found to be the same with those in the vdW system and the mean field theory in nearly all the cases, so they satisfy the scaling laws. Recently, critical exponents beyond mean field theory have been found, which even violate the scaling lawsHu:2024ldp ; Kong:2025lqz . People also found that black holes can undergo Joule-Thomson expansionOkcu:2016tgt , have reentrant phase transitionAltamirano:2013ane and triple pointAltamirano:2013uqa , etc.
Given the thermodynamic similarities between the F(L)RW(Friedmann-(Lemaître)-Robertson-Walker) universe and black holes, one expects the FRW universe to also exhibit a vdW-like equation of state [^2] It should be noted that in the cosmology community, ‘equation of state’ is usually referred to as the relation between pressure $p$ and energy density $\rho$, i.e. $p=\omega\rho$, which is different from the vdW-like equation. In this paper, we use ‘equation of state’ to refer to the vdW-like equation as is often done in the study of black hole phase transitions.. The thermodynamic pressure of the FRW universe is identified with the work densityHayward:1997jp ; Cai:2005ra ; Akbar:2006kj ; Cai:2006rs ; Abdusattar:2021wfv ; Kong:2021dqd ; Kong:2022xny ; Abdusattar:2023hlj ; Chu:2025zuz ,
@@
where $\rho$ and $p$ are the energy density and pressure of the perfect fluid respectively. It can be written as an equation of state $P=P(V,T)$, where $V$ is the thermodynamic volume and $T$ is the Hawking temperature of the FRW universe. The equation of state depends on the theory of gravity and the dimension of spacetime. For the (3+1)-dimensional FRW universe, it has been obtained in Einstein gravityAbdusattar:2021wfv , scalar-tensor theoriesKong:2021dqd ; Abdusattar:2023hlj , the brane world scenarioKong:2022xny , and quasi-topological gravityChu:2025zuz , etc. It has also been obtained for the high dimensional FRW universe in Einstein-Gauss-Bonnet gravitySaavedra:2023lds  [^3] I thank Haximjan Abdusattar for drawing my attention to this paper.. In many cases, the vdW-like phase transitions are found, but they occur above the critical temperature, which is different from the black hole cases.
In this paper, we are interested in the behavior of the perfect fluid in the FRW universe. The perfect fluid in the FRW universe also has a pressure as well as an energy density. If the perfect fluid is in equilibrium with the Hawking radiation from the apparent horizon of the FRW universe, its temperature is also equal to the Hawking temperature. From the corresponding Friedmann’s equations, one can also get the equation of state for the perfect fluid in the FRW universe. The dimension of spacetime also affects the thermodynamic behaviors of the perfect fluid, so we study the perfect fluid in the general $D$-dimensional(or $d+1$ dimensional) FRW universe instead of the $4$-dimensional(or $3+1$ dimensional) FRW universe. In this paper, we only consider three common tensor theories of gravity, i.e. Einstein gravity, Gauss-Bonnet gravity and Lovelock gravityLovelock:1971yv ; Deruelle:1989fj ; Padmanabhan:2013xyr , and obtain the corresponding equation of state for the perfect fluid in the $D$-dimensional FRW universe. In some cases, the vdW-like phase transitions are also found, which depend on the theory of gravity and dimension of the FRW universe as expected. Phase transitions of the perfect fluid also occur above the critical temperature.
@@ II A Brief Introduction of the $D$-dimensional FRW Universe
where $a(t)$ is the time-dependent scale factor [^4] In the following, we may use $a$ instead of $a(t)$ for convenience., $k$ is the spatial curvature (with +1,0,1+1,0,-1 corresponding to spatially closed, flat and open respectively), and $\mathrm{d}\Omega_{n-1}^{2}$ is the metric of the unit (n1)(n-1)-dimensional sphere. One can also rewrite the metric as
@@ III Equation of the Perfect Fluid in Einstein Gravity
The equation of state can be regarded as a function $p=p(R_{A},T)$, in order to show its properties, one can either keep $R_{A}$ or $T$ fixed mathematically. However, for dynamical spacetime like the FRW universe, $R_{A}$ is not fixed, so we only fix $T$ and study the behavior of the isothermal lines, which shows the relation between the pressure and the radius of the apparent horizon. If $n=3$, for the isothermal line with temperature $T$, the pressure has a minimum $p_{min}=-2\pi T^{2}$ at $R=1/(4\pi T)$, and becomes zero at $R=1/(8\pi T)$, see FIG.1.666Here the temperature $T$ is given as 0.11, 0.1, 0.09 without units, which seems strange for some readers. It should be noted that the figures are drawn to show the shape of the isothermal lines and the variation trend of the pressure with the temperature, so the specific value or unit is not so important. It is also a frequently used method in the references about black hole phase transitions, e.g.Kubiznak:2012wp . The values of the temperature in the following figures are chosen on the same spirit.
@@ VI Conclusions and Discussions
The results show that critical points and phase transitions of the perfect fluid depend on both the theories of gravity and the dimensions of the FRW universe. If the critical point exists, phase transitions always take place above the critical temperature, which appears to be a universal feature.